# Learn from Failure: Fine-Tuning LLMs with Trial-and-Error Data for Intuitionistic Propositional Logic Proving

### v1

@@ 5.4 Ablation Study
To evaluate the effectiveness of training with trial-and-error, we craft an ablated version of our method where the LLM is fined-tuned with data of the corrected path only and do inference in the same way as our method (i.e., producing 1 tactic at a time and applying Lean for state checking). We denote the ablated version as Model - proof w/o t.a.e.. For both methods, we mark the search attempt as failed if the tactic induces a Lean error, or the search exceeds the 1500-word limit. The result is shown in the Table 3. The difference between the success rates of the two models is significant, which reaches $29.4\%$. This clearly shows that failed search states and trial-and-error information tremendously enhance the model’s capability to solve theorem-proving tasks.
@@
5.5 Exploratory Study: Training Proof Length Affect Model Performance
Since the FPS algorithm of PropL dataset can generate multiple proofs with variable length, we conduct an exploratory study to assess the impact of proof length on model performance. We fine-tune two models using proofs with different lengths of trial-and-error information. For the first model, which is our TrialMaster, the training data is derived by randomly selecting two out of the shortest four proofs from the ten available proofs for each theorem in PropL. We denote it as Model - short proof w/ t.a.e. In contrast, the training data of the second model is formed by randomly selecting two proofs from the ten available for each theorem, irrespective of their lengths. We denote it as Model - long proof w/ t.a.e. For both models, we use greedy search to let them generate one tactic for each state. We evaluate the models on our 1000 OOD testing set. The results are shown in the Table 4. A higher success rate is observed in the model trained with shorter proofs. This can be attributed to the fact that as the proof with trial-and-error information becomes longer, there is too much trial-and-error information that may detrimentally affect the model’s performance, as too many failed search paths may lower the quality of the training data.
@@ 6 Conclusion and Future Work
In this paper, we study Automated Theorem Proving in formalized environments. We create a complete, scalable, and representative data set of intuitionistic propositional logic theorems in Lean. We demonstrate that leveraging information from failed search states and backtracking not only teaches models how to backtrack effectively but also helps in developing better tactics than those generated by models trained without access to backtracking insights. We release our datasets on GitHub and Huggingface [^1] PropL dataset is available at https://huggingface.co/datasets/KomeijiForce/PropL.
@@
Appendix A Uniformly Distributed Data Explanation
@@ A.1 Catalan Number
A.2 Bijection between Propositions and Natural Numbers
@@
This counting facilitates an efficient ranking of all full binary trees with $n$ internal nodes. This ranking process inherently establishes a bijection with the set of natural numbers, allowing for a clear correspondence between each full binary tree and a unique natural number. Consequently, this sets the stage for a detailed examination of two critical processes: encoding (see Algorithm 1), which involves mapping a proposition tree to a natural number, and decoding (see Algorithm 2), which entails mapping a natural number back to a corresponding proposition tree. These processes are vital for effectively establishing and utilizing the bijection between full binary trees and natural numbers.
@@
Given the uncertainty surrounding the proof lengths when generating propositions, our approach involves uniform sampling for proposition selection. This selection is later refined by excluding propositions according to their proof lengths as computed by the proof generation algorithm. From a probabilistic perspective, this method is the same as pre-excluding propositions exceeding a certain proof length from the entire sampling space prior to the random selection of propositions.
Appendix B Examples
@@
{spverbatim}
variable (p1 p2 p3 p4 p5 : Prop) theorem thm_5_vars_45663772897 : (((p1 ∨ p2) → False) → ((p1 → False) ∧ (p2 → False))) := by – Implications on the right can always be decomposed. – Introduce an assumption h1 that says ((p1 ∨ p2) → False) intro h1 – Now we want to show ((p1 → False) ∧ (p2 → False)) – Conjunctions on the right can always be decomposed. – We then need to show (p1 → False) and (p2 → False) separately. apply And.intro – We are showing (p1 → False). – Implications on the right can always be decomposed. – We introduce assumption h2 for p1. And we try to show False. intro h2 – We want to use the implication h1. So we show its premise. have h3 : (p1 ∨ p2) := by – Show the left disjunct. (The right adjunct leads to an TAE) apply Or.inl – One of the premise coincides with the conclusion. exact h2 – We have shown the premise of h1 (p1 ∨ p2), – we can now drive its conclusion (False), denoted by h4. let h4 := h1 h3 – False on the left can always be used. apply False.elim h4 – We have shown (p1 → False) and now we show (p2 → False). – Implications on the right can always be decomposed. – We introduce assumption h2 for p2. And we try to show False. intro h5 – We want to use the implication h1. So we show its premise. have h6 : (p1 ∨ p2) := by – Show the right disjunct. (The left adjunct leads to an TAE) apply Or.inr – One of the premise coincides with the conclusion. exact h5 – We have shown the premise of h1 (p1 ∨ p2), – we can now drive its conclusion (False), denoted by h7. let h7 := h1 h6 – False on the left can always be used. apply False.elim h7
Figure 6: Lean Proof

### v3

@@ 5.4 Ablation Study
To evaluate the effectiveness of training with trial-and-error, we craft an ablated version of our method where the LLM is fined-tuned with data of the correct path only and do inference in the same way as our method (i.e., producing one tactic at a time and applying Lean for state checking). We denote the ablated version as Model - proof w/o t.a.e.. For both methods, we mark the search attempt as failed if the tactic induces a Lean error, or the search exceeds the 1500-word limit. The result is shown in the Table 3. The difference between the success rates of the two models is $29.4\%$, which is significant. This clearly shows that failed search states and trial-and-error information tremendously enhance the model’s capability to solve theorem-proving tasks.
@@
5.5 Exploratory Study: Effect of Training Proof Length on Model Performance
Since the FPS algorithm of PropL dataset can generate multiple proofs with variable length, we conduct an exploratory study to assess the impact of proof length on model performance. We fine-tune two models using proofs with different lengths of trial-and-error information. For the first model, which is our TrialMaster, the training data is derived by randomly selecting two out of the shortest four proofs from the ten available proofs for each theorem in PropL. We denote it as Model - short proof w/ t.a.e. In contrast, the training data of the second model is formed by randomly selecting two proofs from the ten available for each theorem, irrespective of their lengths. We denote it as Model - long proof w/ t.a.e. For both models, we use greedy search to let them generate one tactic for each state. We evaluate the models on our 1000 OOD test set. The results are shown in the Table 4. A higher success rate is observed in the model trained with shorter proofs. This can be attributed to the fact that as the proof with trial-and-error information becomes longer, there is too much trial-and-error information that may detrimentally affect the model’s performance, as too many failed search paths may lower the quality of the training data.
@@ 6 Conclusion and Future Work
In this paper, we study Automated Theorem Proving in formalized environments. We create a complete, scalable, and representative dataset of intuitionistic propositional logic theorems in Lean. We demonstrate that leveraging information from failed search states and backtracking not only teaches models how to backtrack effectively, but also helps in developing better tactics than those generated by models trained without access to backtracking insights. We release our datasets on GitHub and Huggingface. [^1] PropL dataset is available at https://huggingface.co/datasets/KomeijiForce/PropL. Model weights are available at https://huggingface.co/KomeijiForce/llama-2-7b-propositional-logic-prover. Generation codes for theorems and proofs are available at https://github.com/ucsd-atp/PropL.
@@
Acknowledgments
This work is supported by the National Science Foundation under grants CCF-1955457 and CCF-2220892. This work is also sponsored in part by NSF CAREER Award 2239440, NSF Proto-OKN Award 2333790, as well as generous gifts from Google, Adobe, and Teradata.
Appendix A Inference Rules for Intuitionistic Propositional Logic
Figure 6 shows the inference rules for intuitionistic propositional logic. Rules relate sequents of the form $\Gamma\vdash A$, where $\Gamma$ is an unordered list of assumptions. A derivation is constructed from the axiom (Assumption) using the derivation rules until $\Gamma$ is empty.
For example, the ($\to$-I) rule says that from a derivation of $\Gamma\vdash B$ under the assumption of $\Gamma,A$, we can get a derivation of $\Gamma\vdash A\to B$. The ($\to$-E) rule says that from a derivation of $\Gamma\vdash A\to B$ and a derivation of $\Gamma\vdash A$ we can derive $\Gamma\vdash B$.
[…]
Γ, A ⊢A(Assumption)
[…]
Γ⊢T(T-I)
[…]
Γ⊢FΓ⊢A(F-E)
[…]
Γ⊢A
Γ⊢BΓ⊢A ∧B($\land$-I)
[…]
Γ⊢A ∧BΓ⊢A($\land$-E1)
[…]
Γ⊢A ∧BΓ⊢B($\land$-E2)
[…]
Γ⊢AΓ⊢A ∨B($\lor$-I1)
[…]
Γ⊢BΓ⊢A ∨B($\lor$-I2)
[…]
Γ⊢A ∨B
Γ, A ⊢C
Γ, B ⊢CΓ⊢C($\lor$-E)
[…]
Γ, A ⊢BΓ⊢A →B($\to$-I)
[…]
Γ⊢A →B
Γ⊢A Γ⊢B($\to$-E)
Figure 6: Natural Deduction Rules for Intuitionistic Propositional Logic
Here are an example derivation.
$1 → p$2.
[…]
Appendix B Uniformly Distributed Data Explanation
@@ B.1 Catalan Number
B.2 Bijection between Propositions and Natural Numbers
@@
This counting facilitates an efficient ranking of all full binary trees with $n$ internal nodes. This ranking process inherently establishes a bijection with the set of natural numbers, allowing for a clear correspondence between each full binary tree and a unique natural number. Consequently, this sets the stage for a detailed examination of two critical processes: encoding (see Algorithm 1), which involves mapping a proposition tree to a natural number, and decoding (see Algorithm 2), which entails mapping a natural number back to a corresponding proposition tree.
@@
Given the uncertainty surrounding the proof lengths when generating propositions, our approach involves uniform sampling for proposition selection. This selection is later refined by excluding propositions according to their proof lengths as computed by the proof generation algorithm.
Appendix C Examples
@@
variable (p1 p2 p3 p4 p5 : Prop)
theorem : (((p1 p2) False) ((p1 False) (p2 False))) := by
Implications on the right can always be decomposed.
Introduce an assumption h1 that says ((p1 p2) False)
intro h1
Now we want to show ((p1 False) (p2 False))
Conjunctions on the right can always be decomposed.
We then need to show (p1 False) and (p2 False) separately.
apply And.intro
We are showing (p1 False).
Implications on the right can always be decomposed.
We introduce assumption h2 for p1. And we try to show False.
intro h2
We want to use the implication h1. So we show its premise.
have h3 : (p1 p2) := by
Show the left disjunct. (The right adjunct leads to an TAE)
apply Or.inl
One of the premise coincides with the conclusion.
exact h2
We have shown the premise of h1 (p1 p2),
we can now drive its conclusion (False), denoted by h4.
let h4 := h1 h3
False on the left can always be used.
apply False.elim h4
We have shown (p1 False) and now we show (p2 False).
Implications on the right can always be decomposed.
We introduce assumption h2 for p2. And we try to show False.
intro h5
We want to use the implication h1. So we show its premise.
have h6 : (p1 p2) := by
Show the right disjunct. (The left adjunct leads to an TAE)
apply Or.inr
One of the premise coincides with the conclusion.
exact h5
We have shown the premise of h1 (p1 p2),
we can now drive its conclusion (False), denoted by h7.
let h7 := h1 h6
False on the left can always be used.
apply False.elim h7
Figure 7: An example of an intuitionistic propositional logic theorem with its proof in Lean